# BTL-UI: Blink-Think-Link Reasoning Model for GUI Agent

### v1

@@ 1 Introduction
However, current mainstream systems primarily adopt two approaches. The first relies on supervised fine-tuning (SFT) to align model behavior with task objectives, but this method faces two major limitations: a strong dependence on large-scale expert-labeled data and limited generalization capability when faced with out-of-distribution scenarios. The second approach involves rule-based reinforcement fine-tuning (RFT) guo2025deepseek , as shown in Figure 1 (a), which enhances generalization in complex tasks by using a structured reasoning format: intermediate cognitive steps are encapsulated within ¡think¿ tags, and final decisions are expressed through ¡answer¿ tags. While effective in improving task performance, these methods qin2025ui ; xia2025gui ; liu2025infigui exhibit two critical shortcomings: (1) significant deviation from natural human-GUI interaction patterns, and (2) excessive focus on interaction outcomes while lacking effective process-oriented reward mechanisms.
@@
  • ¡blink¿: Where relevant areas of the screen are rapidly located, analogous to saccadic eye movements. The visual attention-related region-of-interest information is encapsulated within ¡blink¿¡/blink¿ tags.
  • ¡think¿: Where the system engages in high-level reasoning and decision-making, mirroring cognitive task planning. The reasoning processes are recorded in ¡think¿¡/think¿ tags.
  • ¡link¿: Where actionable commands are generated for precise execution, reflecting human action selection mechanisms. The action commands are output in ¡link¿¡/link¿ tags.
@@ 3.2 Blink Data Generation
Figure 3: Two-stage data construction pipeline. In the first stage, the basic properties of UI elements are obtained by a parsing model. To eliminate the redundancy of the number and attributes of elements, the analysis model in the second stage simplifies the list to $\lambda$ elements with their positions (¡bbox¿), while the reserved ¡caption¿ attribute indicates whether the element is interactive. In the example shown in the figure, the instruction for the current step is “Use the GPS to locate a nearby museum and then book a ride with Lyft.” Accordingly, the most relevant element in the Blink output is the “Maps & Navigation” app with ¡ID¿10¡/ID¿.
@@
where $\mathcal{E}_{\text{ROI}}\subseteq\mathcal{E}$ denotes the resulting set of $\lambda$ candidate elements that are most relevant to the current task step during blink phases, encapsulated within ¡blink¿¡/blink¿ tags. The choice of the number of elements after filtering, $\lambda$, achieves a trade-off between BTL performance and efficiency. This attention-guided annotation not only mimics human visual focus during blink phases but also provides a high-quality reference for optimizing the agent policy.
@@ 5 Conclusion and Limitations
We propose the BTL framework, an innovative GUI interaction architecture inspired by the biological cognitive paradigm of ”Blink-Think-Link”. This framework simulates the human closed-loop system of visual perception, cognitive decision-making, and action execution during GUI operations, overcoming the limitations of traditional outcome-driven RFT approaches. Experimental results show that the BTL-UI agent, developed under this framework, achieves significant performance improvements across a variety of GUI interaction tasks.
@@
Limitations. The proposed BTL framework introduces ¡blink¿ tag outputs compared to conventional Think-Answer structured outputs. Although the blink-generated ROI regions are adaptive and can be empty (zero-length), they typically increase the output sequence length in most cases. While demonstrating performance improvements across various GUI task metrics, this design incurs additional processing overhead.
@@ Appendix B Prompt
You are a GUI Agent capable of reasoning based on user instructions, action history, and the current screenshot. You should first observe the layout of the screenshot and extract $N$ elements RELATED TO the user instruction, where 0 ¡= $N$ ¡=5. Next, think about the reasoning process BASED ON the observations and instructions in your mind, and then provide the user with the answer.
The observation process (can be None if $N$ == 0), reasoning process and answer are enclosed within ¡blink¿¡/blink¿, ¡think¿¡/think¿ and ¡link¿¡/link¿ tags, respectively, i.e.,
¡blink¿
¡element¿¡id¿1¡/id¿¡bbox¿[x0, y0, x1, y1]¡/bbox¿¡caption¿dynamic¡/caption¿¡/element¿
¡element¿¡id¿2¡/id¿¡bbox¿[x2, y2, x3, y3]¡/bbox¿¡caption¿static¡/caption¿¡/element¿
¡element¿¡id¿3¡/id¿…..¡/element¿
¡element¿¡id¿4¡/id¿…..¡/element¿
¡/blink¿
¡think¿ reasoning process here ¡/think¿
¡link¿ answer([”Plan”: …, ”Action”: ”function”: …, …]) ¡/link¿.
where captions must be one of [dynamic, static], ”dynamic” refers to the interactive area, and ”static” refers to the non-interactive areas, such as text and diagrams in the screenshot.
And the observation can be ¡blink¿ None ¡/blink¿, if $N$ == 0.

### v2

@@ 1 Introduction
However, current mainstream systems primarily adopt two approaches. The first relies on supervised fine-tuning (SFT) to align model behavior with task objectives, but this method faces two major limitations: a strong dependence on large-scale expert-labeled data and limited generalization capability when faced with out-of-distribution scenarios. The second approach involves rule-based reinforcement fine-tuning (RFT) guo2025deepseek , as shown in Figure 1 (a), which enhances generalization in complex tasks by using a structured reasoning format: intermediate cognitive steps are encapsulated within <think> tags, and final decisions are expressed through <answer> tags. While effective in improving task performance, these methods qin2025ui ; xia2025gui ; liu2025infigui exhibit two critical shortcomings: (1) significant deviation from natural human-GUI interaction patterns, and (2) excessive focus on interaction outcomes while lacking effective process-oriented reward mechanisms.
@@
  • <blink>: Where relevant areas of the screen are rapidly located, analogous to saccadic eye movements. The visual attention-related region-of-interest information is encapsulated within <blink></blink> tags.
  • <think>: Where the system engages in high-level reasoning and decision-making, mirroring cognitive task planning. The reasoning processes are recorded in <think></think> tags.
  • <link>: Where actionable commands are generated for precise execution, reflecting human action selection mechanisms. The action commands are output in <link></link> tags.
@@ 3.2 Blink Data Generation
Figure 3: Two-stage data construction pipeline. In the first stage, the basic properties of UI elements are obtained by a parsing model. To eliminate the redundancy of the number and attributes of elements, the analysis model in the second stage simplifies the list to $\lambda$ elements with their positions (<bbox>), while the reserved <caption> attribute indicates whether the element is interactive. In the example shown in the figure, the instruction for the current step is “Use the GPS to locate a nearby museum and then book a ride with Lyft.” Accordingly, the most relevant element in the Blink output is the “Maps & Navigation” app with <ID>10</ID>.
@@
where $\mathcal{E}_{\text{ROI}}\subseteq\mathcal{E}$ denotes the resulting set of $\lambda$ candidate elements that are most relevant to the current task step during blink phases, encapsulated within <blink></blink> tags. The choice of the number of elements after filtering, $\lambda$, achieves a trade-off between BTL performance and efficiency. This attention-guided annotation not only mimics human visual focus during blink phases but also provides a high-quality reference for optimizing the agent policy.
@@ 5 Conclusion and Limitations
We propose the BTL framework, an innovative GUI interaction architecture inspired by the biological cognitive paradigm of "Blink-Think-Link". This framework simulates the human closed-loop system of visual perception, cognitive decision-making, and action execution during GUI operations, overcoming the limitations of traditional outcome-driven RFT approaches. Experimental results show that the BTL-UI agent, developed under this framework, achieves significant performance improvements across a variety of GUI interaction tasks.
@@
Limitations. The proposed BTL framework introduces <blink> tag outputs compared to conventional Think-Answer structured outputs. Although the blink-generated ROI regions are adaptive and can be empty (zero-length), they typically increase the output sequence length in most cases. While demonstrating performance improvements across various GUI task metrics, this design incurs additional processing overhead.
@@ Appendix B Prompt
You are a GUI Agent capable of reasoning based on user instructions, action history, and the current screenshot. You should first observe the layout of the screenshot and extract $N$ elements RELATED TO the user instruction, where 0 <= $N$ <=5. Next, think about the reasoning process BASED ON the observations and instructions in your mind, and then provide the user with the answer.
The observation process (can be None if $N$ == 0), reasoning process and answer are enclosed within <blink></blink>, <think></think> and <link></link> tags, respectively, i.e.,
<blink>
<element><id>1</id><bbox>[x0, y0, x1, y1]</bbox><caption>dynamic</caption></element>
<element><id>2</id><bbox>[x2, y2, x3, y3]</bbox><caption>static</caption></element>
<element><id>3</id>…..</element>
<element><id>4</id>…..</element>
</blink>
<think> reasoning process here </think>
<link> answer(["Plan": …, "Action": "function": …, …]) </link>.
where captions must be one of [dynamic, static], "dynamic" refers to the interactive area, and "static" refers to the non-interactive areas, such as text and diagrams in the screenshot.
And the observation can be <blink> None </blink>, if $N$ == 0.